2. Pia invests $2500 in an account that earns simple interest. At the end of each month, she earns $11.25 in interest. a What annual rate of simple interest is Pia carning Round your answer to two decimal places. b) How much money will be in her account after 7 years? C. How long will it take for her money to double?


1
Expert's answer
2021-01-21T11:27:25-0500

a) Calculate the interest "I"

"I=Prt"

"r=I/Pt =\frac{11.25}{2500\times 1/12}=0.054=5.4\%"

b)Calculate the total amount after 7 years

"A=P(1+rt)=2500(1+0.054\times 7)=\$3445"

c) Find "t" such that, "A=2P"

"2P=P(1+rt)"

"2=1+rt"

t"=1\/r=1\/0.054=18.52years"